# Breaking the Orthogonality Barrier in Quantum LDPC Codes

### v1

@@ 2 Problem setting: low-weight logical operators induced by check-row removal
$d_{X}^{(\mathrm{res})}$ $:=\min\{\operatorname{wt}(\boldsymbol{x}):\boldsymbol{x}\in\operatorname{Row}(\tilde{H}_{X})\cap C_{Z}\setminus C_{X}^{\perp}\},$ (5)
@@
$d_{Z}^{(\mathrm{res})}$ $:=\min\{\operatorname{wt}(\boldsymbol{z}):\boldsymbol{z}\in\operatorname{Row}(\tilde{H}_{Z})\cap C_{X}\setminus C_{Z}^{\perp}\},$ (7)
@@
The proposed design aims to make these bounds as large as possible. In [12], binary submatrices corresponding to codewords are lifted to a nonbinary field so that they no longer represent codewords; this can be viewed as increasing these residual-based bounds via nonbinary lifting.
@@ 3 Method of matrix design
and $d_{X}^{(\mathrm{res})},d_{Z}^{(\mathrm{res})}$ are large. We define generalized Hagiwara–Imai codes, i.e., mother matrix pairs with block-circulant structure, and develop a general theory for imposing only the active orthogonality $H_{X}H_{Z}^{\mathsf{T}}=0$ between the active matrices $H_{X}$ and $H_{Z}$.
@@ 3.2 Necessary condition for latent non-orthogonality
Assume $\Psi_{r}=0$ for all $r\in[L/2]\setminus\Delta$. For any $i\in[L_{2}]\setminus[J]$ and $k\in[J]$, $r=(k-i)\bmod\,L_{2}$ does not belong to $\Delta$. By assumption $\Psi_{r}=0$, hence for all such $i,k$, $(\hat{H}_{X}(\hat{H}_{Z})^{\mathsf{T}})_{i,k}=\Psi_{(k-i)\bmod\,L_{2}}=0.$ Therefore $H_{X}(\tilde{H}_{Z})^{\mathsf{T}}=0$, and similarly $H_{Z}(\tilde{H}_{X})^{\mathsf{T}}=0$, a contradiction. If all pairs $(i,j)\in[L/2]^{2}\setminus\Delta$ commute, then each term $F_{u}G_{r-u}+G_{r-u}F_{u}$ vanishes over $\mathbb{F}_{2}$, so $\Psi_{r}=0$ for all $r\notin\Delta$. ∎
@@ 6.1 Structure and active orthogonality
$\begin{aligned} \hat{H}_{X}=&\left(\begin{array}[]{llllll|llllll}F_{0}&F_{1}&F_{2}&F_{3}&F_{4}&F_{5}&G_{0}&G_{1}&G_{2}&G_{3}&G_{4}&G_{5}\\ F_{5}&F_{0}&F_{1}&F_{2}&F_{3}&F_{4}&G_{5}&G_{0}&G_{1}&G_{2}&G_{3}&G_{4}\\ F_{4}&F_{5}&F_{0}&F_{1}&F_{2}&F_{3}&G_{4}&G_{5}&G_{0}&G_{1}&G_{2}&G_{3}\\ \hline\cr F_{3}&F_{4}&F_{5}&F_{0}&F_{1}&F_{2}&G_{3}&G_{4}&G_{5}&G_{0}&G_{1}&G_{2}\\ F_{2}&F_{3}&F_{4}&F_{5}&F_{0}&F_{1}&G_{2}&G_{3}&G_{4}&G_{5}&G_{0}&G_{1}\\ F_{1}&F_{2}&F_{3}&F_{4}&F_{5}&F_{0}&G_{1}&G_{2}&G_{3}&G_{4}&G_{5}&G_{0}\end{array}\right).\\ \hat{H}_{Z}=&\left(\begin{array}[]{llllll|llllll}G_{0}^{\prime}&G_{1}^{\prime}&G_{2}^{\prime}&G_{3}^{\prime}&G_{4}^{\prime}&G_{5}^{\prime}&F_{0}^{\prime}&F_{1}^{\prime}&F_{2}^{\prime}&F_{3}^{\prime}&F_{4}^{\prime}&F_{5}^{\prime}\\ G_{5}^{\prime}&G_{0}^{\prime}&G_{1}^{\prime}&G_{2}^{\prime}&G_{3}^{\prime}&G_{4}^{\prime}&F_{5}^{\prime}&F_{0}^{\prime}&F_{1}^{\prime}&F_{2}^{\prime}&F_{3}^{\prime}&F_{4}^{\prime}\\ G_{4}^{\prime}&G_{5}^{\prime}&G_{0}^{\prime}&G_{1}^{\prime}&G_{2}^{\prime}&G_{3}^{\prime}&F_{4}^{\prime}&F_{5}^{\prime}&F_{0}^{\prime}&F_{1}^{\prime}&F_{2}^{\prime}&F_{3}^{\prime}\\ \hline\cr G_{3}^{\prime}&G_{4}^{\prime}&G_{5}^{\prime}&G_{0}^{\prime}&G_{1}^{\prime}&G_{2}^{\prime}&F_{3}^{\prime}&F_{4}^{\prime}&F_{5}^{\prime}&F_{0}^{\prime}&F_{1}^{\prime}&F_{2}^{\prime}\\ G_{2}^{\prime}&G_{3}^{\prime}&G_{4}^{\prime}&G_{5}^{\prime}&G_{0}^{\prime}&G_{1}^{\prime}&F_{2}^{\prime}&F_{3}^{\prime}&F_{4}^{\prime}&F_{5}^{\prime}&F_{0}^{\prime}&F_{1}^{\prime}\\ G_{1}^{\prime}&G_{2}^{\prime}&G_{3}^{\prime}&G_{4}^{\prime}&G_{5}^{\prime}&G_{0}^{\prime}&F_{1}^{\prime}&F_{2}^{\prime}&F_{3}^{\prime}&F_{4}^{\prime}&F_{5}^{\prime}&F_{0}^{\prime}\end{array}\right).\end{aligned}$ (15)
Here, for simplicity, $A^{\prime}$ denotes the transpose of $A$. With $\Delta=\{0,1,2,4,5\}$,
@@ 6.2 Design goal and commutation table
As a schematic commutation table, suppose all pairs $(F_{i},G_{j})$ commute except $(i,j)=(0,3)$. The table below shades commuting pairs in $\Gamma$ with 1:
@@ 6.5 Minimum-distance evaluation
Thus $\Psi_{3}$ is a sum of four permutation matrices with disjoint supports, so both row and column weights are 4. Therefore any codeword $\boldsymbol{u}$ of $H_{Z}(\tilde{H}_{X})^{\mathsf{T}}$ decomposes as $\boldsymbol{u}=(\boldsymbol{u}_{3},\boldsymbol{u}_{4},\boldsymbol{u}_{5})$ where each $\boldsymbol{u}_{i}$ is a codeword of $\mathrm{Ker}(\Psi_{3}^{\mathsf{T}})$. From the structure of $\mathrm{Ker}(\Psi_{3}^{\mathsf{T}})$ and $\mathrm{Ker}(\Psi_{3})$, the support splits and weights are multiples of 4. The minimum distance of $\mathrm{Ker}(\Psi_{3}^{\mathsf{T}})$ is 4. We confirmed the existence of a weight-48 $\boldsymbol{x}$ constructed from a weight-4 codeword of $\Psi_{3}^{\mathsf{T}}$. Therefore $d_{X}^{(\mathrm{res})},d_{Z}^{(\mathrm{res})}\leq 48$, and consequently $d_{\min}\leq 48$. We do not prove a matching lower bound. Randomized searches did not find smaller logical errors, and large-scale decoding experiments did not exhibit logical failures, so we conjecture that this upper bound is tight.

### v2

@@ 2 Problem setting: low-weight logical operators induced by check-row removal
$d_{X}^{(\mathrm{lat})}$ $:=\min\{\operatorname{wt}(\boldsymbol{x}):\boldsymbol{x}\in\operatorname{Row}(\tilde{H}_{X})\cap C_{Z}\setminus C_{X}^{\perp}\},$ (5)
@@
$d_{Z}^{(\mathrm{lat})}$ $:=\min\{\operatorname{wt}(\boldsymbol{z}):\boldsymbol{z}\in\operatorname{Row}(\tilde{H}_{Z})\cap C_{X}\setminus C_{Z}^{\perp}\},$ (7)
@@
The proposed design aims to make these bounds as large as possible. In [12], binary submatrices corresponding to codewords are lifted to a nonbinary field so that they no longer represent codewords; this can be viewed as increasing these latent-based bounds via nonbinary lifting.
@@ 3 Method of matrix design
and $d_{X}^{(\mathrm{lat})},d_{Z}^{(\mathrm{lat})}$ are large. We define generalized Hagiwara–Imai codes, i.e., mother matrix pairs with block-circulant structure, and develop a general theory for imposing only the active orthogonality $H_{X}H_{Z}^{\mathsf{T}}=0$ between the active matrices $H_{X}$ and $H_{Z}$.
@@ 3.2 Necessary condition for latent non-orthogonality
Assume $\Psi_{r}=0$ for all $r\in[L/2]\setminus\Delta$. For any $i\in[L/2]\setminus[J]$ and $k\in[J]$, $r=(k-i)\bmod\,L/2$ does not belong to $\Delta$. By assumption $\Psi_{r}=0$, hence for all such $i,k$, $(\hat{H}_{X}(\hat{H}_{Z})^{\mathsf{T}})_{i,k}=\Psi_{(k-i)\bmod\,L/2}=0.$ Therefore $H_{X}(\tilde{H}_{Z})^{\mathsf{T}}=0$, and similarly $H_{Z}(\tilde{H}_{X})^{\mathsf{T}}=0$, a contradiction. If all pairs $(i,j)\in[L/2]^{2}\setminus\Delta$ commute, then each term $F_{u}G_{r-u}+G_{r-u}F_{u}$ vanishes over $\mathbb{F}_{2}$, so $\Psi_{r}=0$ for all $r\notin\Delta$. ∎
@@ 6.1 Structure and active orthogonality
$\begin{aligned} \hat{H}_{X}=&\left(\begin{array}[]{llllll|llllll}F_{0}&F_{1}&F_{2}&F_{3}&F_{4}&F_{5}&G_{0}&G_{1}&G_{2}&G_{3}&G_{4}&G_{5}\\ F_{5}&F_{0}&F_{1}&F_{2}&F_{3}&F_{4}&G_{5}&G_{0}&G_{1}&G_{2}&G_{3}&G_{4}\\ F_{4}&F_{5}&F_{0}&F_{1}&F_{2}&F_{3}&G_{4}&G_{5}&G_{0}&G_{1}&G_{2}&G_{3}\\ \hline\cr F_{3}&F_{4}&F_{5}&F_{0}&F_{1}&F_{2}&G_{3}&G_{4}&G_{5}&G_{0}&G_{1}&G_{2}\\ F_{2}&F_{3}&F_{4}&F_{5}&F_{0}&F_{1}&G_{2}&G_{3}&G_{4}&G_{5}&G_{0}&G_{1}\\ F_{1}&F_{2}&F_{3}&F_{4}&F_{5}&F_{0}&G_{1}&G_{2}&G_{3}&G_{4}&G_{5}&G_{0}\end{array}\right).\\ \hat{H}_{Z}=&\left(\begin{array}[]{llllll|llllll}G_{0}^{\prime}&G_{5}^{\prime}&G_{4}^{\prime}&G_{3}^{\prime}&G_{2}^{\prime}&G_{1}^{\prime}&F_{0}^{\prime}&F_{5}^{\prime}&F_{4}^{\prime}&F_{3}^{\prime}&F_{2}^{\prime}&F_{1}^{\prime}\\ G_{1}^{\prime}&G_{0}^{\prime}&G_{5}^{\prime}&G_{4}^{\prime}&G_{3}^{\prime}&G_{2}^{\prime}&F_{1}^{\prime}&F_{0}^{\prime}&F_{5}^{\prime}&F_{4}^{\prime}&F_{3}^{\prime}&F_{2}^{\prime}\\ G_{2}^{\prime}&G_{1}^{\prime}&G_{0}^{\prime}&G_{5}^{\prime}&G_{4}^{\prime}&G_{3}^{\prime}&F_{2}^{\prime}&F_{1}^{\prime}&F_{0}^{\prime}&F_{5}^{\prime}&F_{4}^{\prime}&F_{3}^{\prime}\\ \hline\cr G_{3}^{\prime}&G_{2}^{\prime}&G_{1}^{\prime}&G_{0}^{\prime}&G_{5}^{\prime}&G_{4}^{\prime}&F_{3}^{\prime}&F_{2}^{\prime}&F_{1}^{\prime}&F_{0}^{\prime}&F_{5}^{\prime}&F_{4}^{\prime}\\ G_{4}^{\prime}&G_{3}^{\prime}&G_{2}^{\prime}&G_{1}^{\prime}&G_{0}^{\prime}&G_{5}^{\prime}&F_{4}^{\prime}&F_{3}^{\prime}&F_{2}^{\prime}&F_{1}^{\prime}&F_{0}^{\prime}&F_{5}^{\prime}\\ G_{5}^{\prime}&G_{4}^{\prime}&G_{3}^{\prime}&G_{2}^{\prime}&G_{1}^{\prime}&G_{0}^{\prime}&F_{5}^{\prime}&F_{4}^{\prime}&F_{3}^{\prime}&F_{2}^{\prime}&F_{1}^{\prime}&F_{0}^{\prime}\end{array}\right).\end{aligned}$ (15)
Here, for simplicity, $A^{\prime}$ denotes the transpose of $A$. For example, the $(0,1)$ block of $\hat{H}_{X}(\hat{H}_{Z})^{\mathsf{T}}$ is
$(\hat{H}_{X}(\hat{H}_{Z})^{\mathsf{T}})_{0,1}$ $=F_{0}G_{1}+F_{1}G_{0}+F_{2}G_{5}+F_{3}G_{4}+F_{4}G_{3}+F_{5}G_{2}$
$\quad+G_{0}F_{1}+G_{1}F_{0}+G_{2}F_{5}+G_{3}F_{4}+G_{4}F_{3}+G_{5}F_{2}$
$=\Psi_{1}.$
With $\Delta=\{0,1,2,4,5\}$,
@@ 6.2 Design goal and commutation table
As a schematic commutation table, suppose all pairs $(F_{i},G_{j})$ commute except $(i,j)=(0,3),(1,2)$. The table below shades commuting pairs in $\Gamma$ with 1:
@@ 6.5 Minimum-distance evaluation
Thus $\Psi_{3}$ is a sum of four permutation matrices with disjoint supports, so both row and column weights are 4. Therefore any codeword $\boldsymbol{u}$ of $H_{Z}(\tilde{H}_{X})^{\mathsf{T}}$ decomposes as $\boldsymbol{u}=(\boldsymbol{u}_{3},\boldsymbol{u}_{4},\boldsymbol{u}_{5})$ where each $\boldsymbol{u}_{i}$ is a codeword of $\mathrm{Ker}(\Psi_{3}^{\mathsf{T}})$. From the structure of $\mathrm{Ker}(\Psi_{3}^{\mathsf{T}})$ and $\mathrm{Ker}(\Psi_{3})$, the support splits and weights are multiples of 4. The minimum distance of $\mathrm{Ker}(\Psi_{3}^{\mathsf{T}})$ is 4. We confirmed the existence of a weight-48 $\boldsymbol{x}$ constructed from a weight-4 codeword of $\Psi_{3}^{\mathsf{T}}$. Therefore $d_{X}^{(\mathrm{lat})},d_{Z}^{(\mathrm{lat})}\leq 48$, and consequently $d_{\min}\leq 48$. We do not prove a matching lower bound. Randomized searches did not find smaller logical errors, and large-scale decoding experiments did not exhibit logical failures, so we conjecture that this upper bound is tight.